# Holomorphic Yang-Mills fields on $B$-branes

### v1

@@ 3.4. The homotopy category
Given $\xi\in{\rm Hom}_{K^{b}_{\rm coh}(Y)}\big{(}{\mathscr{F}}^{\bullet},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{F}}^{\bullet}\big{)}$, according to Remark 7, it determines $\zeta^{i}\in{\rm Hom}\big{(}{\mathscr{H}}^{i},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{H}}^{i}\big{)}.$ As $\nabla^{\bullet}$ is, by hypothesis, a Yang-Mills field then (3.32) is satisfied.
@@
for any variation $\tau$.
The morphism $\tau$ can be extended
@@
$0=\sum_{i}(-1)^{i}\langle K_{\vartheta^{i}},\,\vartheta^{i}(\zeta^{i})\rangle=(-1)^{j}\langle K_{{\vartheta}^{j}},\,\vartheta^{j}(\zeta^{j})\rangle=(-1)^{j}\langle K_{{\vartheta}^{j}},\,\vartheta^{j}(\tau)\rangle.$
@@
From Theorem 6 together with Proposition 14, it follows the corollary.
Corollary 4.
Let $({\mathscr{F}}^{\bullet},\,\delta^{\bullet})$ be a Hermitian $B$-brane on the Hodge manifold $Y,$ such that sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ are reflexive. A gauge field $\psi$ on the brane is a Yang-Mills field iff $\vartheta^{i}$ is a Yang-Mills field on ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ for all $i$.

### v2

@@ 3.2.2. The Yang-Mills functional
Given $\xi\in{\rm Hom}_{D^{b}(X)}\big{(}{\mathscr{F}}^{\bullet},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{F}}^{\bullet}\big{)}$, according to Remark 6, it determines $\xi^{i}\in{\rm Hom}\big{(}{\mathscr{H}}^{i},\,\Omega^{1}\otimes_{\mathscr{O}}{\mathscr{H}}^{i}\big{)}.$ As $\nabla^{\bullet}$ is, by hypothesis, a Yang-Mills field then (3.20) is satisfied.
@@
for any variation $\tau$. The morphism $\tau$ can be extended to a morphism
@@
$0=\sum_{i}(-1)^{i}\langle K_{\vartheta^{i}},\,\vartheta^{i}(\xi^{i})\rangle=(-1)^{j}\langle K_{{\vartheta}^{j}},\,\vartheta^{j}(\xi^{j})\rangle=(-1)^{j}\langle K_{{\vartheta}^{j}},\,\vartheta^{j}(\tau)\rangle.$
@@
4. Final Remarks
The following are some aspects and assumptions made in the article that could be considered to generalize or extend the derived results.
Remark 7.
In the definition of the Yang-Mills functional on the gauge fields on brane ${\mathscr{F}}^{\bullet}$ we assumed that the cohomology sheaves ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ are reflexive. It would be interesting to extend the definition to holomorphic gauge fields on an arbitrary $B$-brane.
Remark 8.
The semisimplicity assumption in Theorem 5 is a strong hypothesis that would be desirable to weaken.
Remark 9.
In Section 2.3.2, we considered the variety of complete flags in ${\mathbb{C}}^{3}$ and proved Theorem 2, which provides an upper bound for the number of gauge fields on this variety. The restriction to this flag variety arises from the fact that, in this case, any Schubert subvariety $Z$ is smooth. Thus, from the vanishing of the first Betti number, we deduce that $H^{1,0}(Z)=0,$ which is a necessary condition for Proposition 6 to hold. Following an alternative approach, could it be possible to prove a version of Theorem 2 for any complete flag variety?
Remark 10.
The definitions introduced in the article and the basic results pertain to $B$-branes over a complex manifold. One may wonder which parts of the article’s content are generalizable to the context of branes over varieties or schemes. It is expected that anything admitting a formulation in categorical language, such as the definition of jet sheaf or Proposition 1, can be translated into the framework of scheme theory. However, the translation of other concepts originating in differential geometry, such as the Yang-Mills functional, is not straightforward.